OpenStudy (anonymous):

Which function below is the inverse of f(x) = The quantity of seven x plus one, over two.? f−1(x) = two over the quantity of seven x plus one f−1(x) = two over the quantity of one minus seven x f−1(x) = The quantity of two x plus one, over seven f−1(x) = The quantity of two x minus one, over seven

OpenStudy (debbieg):

Don't worry about the answer choices. Don't need 'em, you just need to know how to find the inverse. You have: \[\Large f(x)=\frac{ 7x+1 }{2 }\] You want to do what is called "switch and solve" method. First, write the function in y=.... notation (instead of f(x)=): \[\Large y=\frac{ 7x+1 }{2 }\] Now, the "switch". Anywhere that you have a y, make it an x; and anywhere that you have an x, make it a y: \[\Large x=\frac{ 7y+1 }{2 }\] Now, take THAT equation and SOLVE it for y (hence, the "solve" part). That's your inverse function, \(\Large f^{-1}(x)\)

OpenStudy (anonymous):

im confuse, how do i solve for y?

OpenStudy (debbieg):

What class is this? if you are in a class where you are finding inverses of function (probably algebra 2?) then you should know how to solve an equation for a specified variable..... Solve it for y, by isolating the y. Use the properties of equality to "shuffle" the equation around, until you have y=......

OpenStudy (anonymous):

i have algebra 2 i just started. i know i have to isolate y but what bout x? is just confusing

OpenStudy (debbieg):

I'll get you started: we need to get the y alone, so lets multiply both sides by 2: \[\Large 2x=\left(\dfrac{ 7y+1 }{2 } \right)\cdot2\] \[\Large 2x=\left(\dfrac{ 7y+1 }{\cancel 2 } \right)\cdot\cancel 2\] \[\Large 2x= 7y+1 \] what's next?

OpenStudy (anonymous):

-1 both sides?

OpenStudy (debbieg):

I don't understand what you mean, "i know i have to isolate y but what bout x? " What about x?? Isolate the y on one side, and there will be an x on the other side....

OpenStudy (debbieg):

right - good.

OpenStudy (debbieg):

and then just one more step

OpenStudy (anonymous):

2x - 1 = 7y then divide 7 which would be 7x/7 - 1 = y?

OpenStudy (debbieg):

\(\Large 2x-1= 7y\)

OpenStudy (debbieg):

Uhm, yes, divide by 7.... but what you get doesn't look like what you have there??

OpenStudy (anonymous):

2x*

OpenStudy (anonymous):

2x/7 - 1 = y

OpenStudy (debbieg):

NO.... not: \(\Large \dfrac{2x}{7}-1= y\) Look at what you have to do: you need to divide ALL of the LEFT SIDE by 7. Not just the first term.

OpenStudy (anonymous):

ohhh my bad my bdad. 2x/7 - 1/7 = y ?

OpenStudy (debbieg):

\(\Large \dfrac{1}{7}\cdot(2x-1)= 7y\cdot \dfrac{1}{7}\)

OpenStudy (debbieg):

Right! Although I would just put the left side in one rational expression, but what you wrote is equivalent. \[\Large \dfrac{2x-1}{7}= y\]

OpenStudy (anonymous):

thank you very much!

OpenStudy (debbieg):

you're welcome. :) happy to help :)

OpenStudy (anonymous):

can you help me with this one real quick please? A function is created to represent the costs of living per person in the family. What restrictions would be made to the domain? The domain would only include integers. The domain would only include positive integers. The domain would only include positive numbers. The domain would include all real numbers.

OpenStudy (debbieg):

The domain is all the values that x (or the independent variable, the "input") is "allowed" to be. Here, the function is cost of living PER PERSON, so I would assume that the x is the number of people in the family. What KIND of number can that variable be? Can it be negative? can it be something that's NOT an integer (e.g., 1/2)?

OpenStudy (anonymous):

cant be negative and it cant be sumthing thats not an integer. it would be only positive integers? becase there cant be a negative person in the family or half a person in the family?

OpenStudy (anonymous):

@DebbieG

OpenStudy (debbieg):

Exactly. :)

OpenStudy (anonymous):

Which function below is the inverse of f(x) = The quantity of four x minus three, over two. @debbieg

OpenStudy (anonymous):

The domain would include all real numbers. so the last one